OpenStudy (anonymous):

y varies inversely with x. Find the inverse variation equation if x = -2.5 when y = 50.

OpenStudy (shayaan_mustafa):

Hi kena :) How are you dear? Welocme to Openstudy

OpenStudy (anonymous):

lol , who are you ?

OpenStudy (shayaan_mustafa):

I am a user of Openstudy :)

OpenStudy (anonymous):

ohh hey are you good at alg1?

OpenStudy (shayaan_mustafa):

\[\large y\space\alpha\space\frac{1}{x}\] \[\large y=k(\frac{1}{x})\] Put x and y and find k

OpenStudy (shayaan_mustafa):

!!!!

OpenStudy (anonymous):

huh?

OpenStudy (shayaan_mustafa):

I told you to find the k form the given equation dear

OpenStudy (anonymous):

Can you solve the whole thing for me please???

OpenStudy (shayaan_mustafa):

OK. You have given x and y in the question. So put them in the equation like this. \[\large 50=k(\frac{1}{-2.5})\] \[\large -125=k\] So your answer is k=-125

OpenStudy (anonymous):

The answers that is in the multiple choice is A.y/x = -125 B.xy = -125 C.y/x = -20 D.xy = -20

OpenStudy (shayaan_mustafa):

It is B.

OpenStudy (anonymous):

The amount of information retained varies inversely with the number of hours passed since the information was given. If Diana can retain 20 new vocabulary words 1/4 hour after she learns them, how many will she retain 2.5 hours after she reads them?